OpenStudy (anonymous):

Simplify. –15 – (–3) + 4 A.–8 B.–16 C.8 D.–22

OpenStudy (mathstudent55):

-(-) = +

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

thats what i thoght too @zak92997 i just wanted to make sure! thanks!

OpenStudy (mathstudent55):

Do it one step at a time. –15 – (–3) + 4 = -15 + 3 + 4 = -12 + 4 = -8

OpenStudy (anonymous):

no prob